# Planet formation regulated by galactic-scale interstellar turbulence

### v1

@@ 0.5 Properties of local star forming regions (step 1):
In the first instance, we wish to quantify the properties of giant molecular clouds (GMCs) that undergo gravitational collapse to form stars. This can be reframed in terms of quantifying the maximum turbulent size-scale on which a given volume of space in the mid-plane of the galaxy is marginally gravitationally unstable at any given time. All scales smaller than this size-scale will also be bound to this cloud, and therefore this can be considered a unit of star formation, or a star forming region. The problem of quantifying the mass or radius distribution of collapsing star forming regions can be solved via excursion set theory, as has been successfully applied to dark matter halos.[38]
@@
We follow Hopkins (2012),[39] who applied the excursion set formalism in the context of GMCs in the galactic disc. The approach can be summarised as follows, but we refer the interested reader to that work for greater detail. Most generally, it is clear that the density of gas on any given scale can be described as the sum of the contributions to this density at all larger scales. Supersonic turbulence yields a log-normal probability distribution in volume space for the gas density, with a dispersion that is dependent on the local Mach number.[71] Therefore, dispersion in the contribution of any range of scales to the deviation from the median depends on the turbulent velocity on that scale. In a supersonic, rapidly cooling medium, the cascade of energy from large scales to small scales produces an energy spectrum with power-law exponent $p=2$, or turbulent velocity that scales with the square-root of the size-scale.[72] The critical density required for collapse at any intermediate size-scale is then related to the effective pressure support afforded to the medium from the turbulent motions on that scale, via the dispersion relation for a disc of a finite thickness.[73] On the largest scales, this pressure support effectively comes from galactic shear, while on the smallest scales, turbulent fragmentation can be suppressed by the thermal pressure.[74] In principle, the influence of magnetic fields can also be accounted for in this context, but Hopkins (2012) notes that this modification practically amounts to changing the sound speed, with only minor quantitative consequences for our model. In the context of the BHL accretion rate that we eventually compute, modifications of only a factor of order unity are expected due to local magnetic fields.[75, 76] Given the turbulent energy spectrum, a ‘trajectory’ in logarithmic density space can be instantaneously defined as the cumulative sum of perturbations from the largest to smallest scales. For any trajectory, a gravitationally unstable GMC, or star forming region, is defined on the largest length scale which is unstable (if any), all smaller length scales being subsumed within the larger volume.
@@
In our case, we follow similar parameter choices to those adopted by Hopkins (2012). The energy spectrum normalisation is defined such that the velocity dispersion at the scale height $h$ of the gaseous galactic disc $\sigma_{\rm g}(h)=h\Omega=6$ km s-1, where $\Omega=2.6\times 10^{-2}$ Myr-1. We assume a surface density of the galactic disc in the solar neighbourhood $\Sigma=12\,M_{\odot}$ pc-2, yielding a Toomre $Q=1.33$. We always assume a sound speed $c_{\mathrm{s}}=0.2$ km s-1 and define a grid with a maximum scale $R_{\mathrm{max}}=5h$ down to a minimum scale $R_{\mathrm{min}}<0.01$ pc, each grid cell being $95$ percent of the spatial scale of the previous cell.
@@
where $t_{\rm cross}$ is the crossing time-scale and $\eta\approx 1$.[77] An appropriate time-scale ${\tau}_{\mathrm{t,0}}$ over which the trajectory ‘resets’ corresponds to the turbulent time-scale on the scales that dominate density variations. We therefore estimate ${\tau}_{\mathrm{t,0}}\approx\tau_{\mathrm{t}}(h)$. We can then estimate the differential number of star forming region younger than a certain age $\Delta\tau$ within a given volume $\Delta V$:
@@
where $N_{\mathrm{coll}}$ is the number of unstable trajectories from a sample of $N_{\mathrm{traj}}$. We include a factor $1/2^{3}$ since given a region of radius $R_{\mathrm{SFR}}$, any other region with $R_{\mathrm{SFR}}$ is forbidden inside a distance of $2R_{\mathrm{SFR}}$ (such a region would merge with the original region). In the averaging for equation 4 we have ignored an obvious consequence of the random walk. That is that the unstable regions should be spatially correlated, since clouds preferentially collapse in regions of high density at spatial scales larger than the nascent star forming region. However, Hopkins (2012) find that there is little bias for the low mass star forming regions, with $R_{\mathrm{SFR}}\ll h$. These are the regions relevant for our analysis.
We show the outcome of this calculation for $10^{5}$ independent trajectories in Figure Extended Data Figure 1. We plot the number of star forming regions above a given mass $M$ within a spherical volume with radius $R_{\mathrm{obs}}=200$ pc younger than $10$ Myr. This is approximately the radius within which the majority of protoplanetary disc statistics have been compiled,[42] and the dispersal time-scale for protoplanetary discs.[1, 78] We define the mass threshold below which we expect $>10$ star forming regions to be the threshold for being well sampled. This yields a mass limit $M_{\mathrm{local}}\approx 500\,M_{\odot}$. Note, our definition of a star forming region is a region that is collapsing to form stars. This is in contrast to a broader complex that may not be collapsing globally. As an illustrative example, the Taurus-Auriga association has a total mass much larger than $500\,M_{\odot}$.[68] However, the complex as a whole is not collapsing to form stars. Instead collapse and star formation has occurred in several smaller sub-regions. It is these regions, or ‘NESTS’,[79] that are the individual star forming regions in the sense appropriate to this work. In Taurus, these regions are typically a few tens of solar masses. The older region Upper Scorpius has a total stellar mass $\sim 1400\,M_{\odot}$,[80] but is also composed of several individual populations.[81] The nearest individual star forming region with total mass of several $10^{3}\,M_{\odot}$ is the Orion Nebula cluster,[82] at a distance of $\sim 400$ pc.[83] We conclude that, although our approach is approximate in that it neglects some physical processes that may influence local star formation, such as possible nearby supernovae,[84] the upper mass limit we infer is reasonable. Our subsequent results do not depend strongly on this limit, as long as we exclude much more massive regions that may be dispersed prematurely by stellar feedback.
@@
0.6 Collapse of the giant molecular cloud (step 2):
@@
In the early stage after a star forms, it occupies the local star forming region that is undergoing gravitational collapse. Since we are interested in the local density and velocity evolution throughout a star’s lifetime, we require a simple model to follow the collapse of the GMC and formation of stars. Once the GMC becomes gravitationally unstable, it collapses on a free-fall time-scale. It does not do so monolithically, but because smaller sub-regions of the cloud may be higher density (and therefore have shorter free-fall time-scales), they fragment first and continue to accrete from their parent structures.[85] Simultaneously, the cloud is exposed to heating by feedback processes from the stars as they form. A wide range of analytic and semi-analytic approaches have been developed to describe this process.[86] We are helped in this context in our interest in low mass star forming regions that are locally well sampled. Since we do not expect many massive stars or supernovae among these local regions, we will hereafter ignore feedback processes and exclusively consider gravitational collapse.
@@
We adopt the approach of Girichidis et al. (2014)[40] in modelling the rate of star formation during the collapse of the GMC. The approach of those authors is attractive for its simplicity, while also retaining a self-consistent treatment of the gas density distribution as it evolves under gravity. We briefly review the approach here, referring the reader to the original work for details.
@@
We consider a spherical region undergoing free-fall collapse from rest, which may be contained within a larger gravitationally unstable cloud. Girichidis (2014) wrote an invertible, analytic approximation for the density of such a region as a function of time. Given an analytically integrable initial density distribution, there is a corresponding analytic expression for the fraction of an initial gas reservoir that collapses beyond any given threshold as a function of time. The initial density distribution in our collapsing region is log-normal, with a dispersion determined by the Mach number on the scale of collapse. We choose the critical density to be such that the local free-fall time is shorter than $0.1$ Myr. Given the accelerating nature of gravitational collapse, the exact threshold time is irrelevant so long as it is substantially shorter than the time-scale for the global collapse of the star forming region. The star formation rate is defined to be the rate at which material crosses this threshold density. We assume the cloud is instantly dispersed on cessation of the star formation phase, which continues until a star formation efficiency $\epsilon=0.5$ is reached, this value being commensurate with simulations of stellar core formation.[41] We expect stellar feedback, which we do not consider and may shut off star formation in massive clouds, to have a negligible influence on the low mass star forming regions that are our focus. High star formation efficiencies similar to those we assume are also reached for more sophisticated models of star formation in a collapsing cloud for low mass clouds.[86]
0.7 Stellar dynamics in the bound cloud (step 3):
@@
Having defined a local star formation rate, we can draw formation times for an initial stellar population. Since these stars are initially bound to the cloud from which they form, we integrate the orbit of the star within the gravitational potential of the cloud, ignoring interactions with other stars. For convenience we will consider a Plummer potential with a core radius $a=R_{\mathrm{SFR}}/2.6$, such that the outer radius of the star forming region is twice the half-mass radius. Stars are drawn from within the core radius, and we draw an isotropic velocity distribution with a dispersion assuming instantaneous virial equilibrium with the cloud. To track local density and relative velocity of the gas, we assume a static, smooth Plummer density profile.[36] We then evolve the star under the potential, varying the core radius as the spherical star forming region contracts until the final dispersal of the cloud (when $M_{*}=\epsilon M_{\mathrm{tot}}$).
0.8 Evolution of the turbulent medium (step 4):
@@
0.9 Accretion rate onto the star-disc system (step 5):
Bondi-Hoyle-Lyttleton (BHL) accretion has been the topic of many works.[34] In the purest case, it is the gravitational focusing of the test particles (representing the ISM), resulting in a stagnation point in the wake of the star within which material is accreted. The instantaneous Bondi-Hoyle-Lyttleton (BHL) radius within which material is captured is then defined:
@@ 0.9 Accretion rate onto the star-disc system (step 5):
Streamlines with an impact parameter at the BHL radius $R_{\mathrm{BHL}}$ hit a stagnation point along the axis coincident with the relative upstream gas velocity vector. In the ideal case, in an homogeneous medium, they then accrete directly onto the star.
If the tidal radius $R_{\mathrm{T}}<R_{\mathrm{BHL}}$, then this radius limits the material that can be captured by the star. If a star is bound to a mass $M_{\mathrm{b}}$ at a radius $r_{\mathrm{b}}$, then:
@@
Initially, this radius is determined by the collapsing cloud, where $r_{\mathrm{b}}$ is the instantaneous location of the star with respect to the centre of the region. Once the cloud has dispersed, we take the galactic values $r_{\mathrm{b}}=8$ kpc and $M_{\mathrm{b}}=10^{12}\,M_{\odot}$. The accretion radius also cannot exceed the turbulent size scale under consideration, $R_{\mathrm{t}}$. Thus the accretion radius can be defined:
@@
Material within the cross-section carved out by this radius is captured by the star. At any turbulent scale, we can estimate the instantaneous rate at which a star captures material in such a scenario:[33, 87]
@@
We show the outcome for the evolution of the local density, relative velocity and BHL accretion rate (normalised to show the accretion rate assuming a solar mass star) for a subset of six stars in Figure 1. Given that density in particular can vary over several orders of magnitude, we also draw an approximate comparison to the typical densities in observed star forming regions. To do so, we adopt averaged density estimates from Table 4 of Goldsmith et al. (2008)[68] for a number of high density areas within the Taurus star forming complex. We convert the mass estimates $M_{\mathrm{enc}}$ within the quoted surface area $A_{\mathrm{surf}}$ to a density estimate by approximating spherical geometry, such that the volume averaged density is:
@@
0.10 Disc properties (step 6):
BHL accretion in a turbulent medium is unlikely to result in accretion directly onto the star. In an homogeneous medium the velocities of steamlines are zero at the stagnation point. Captured material therefore has zero angular momentum and accretes directly onto the star. However, in an inhomogeneous medium, vorticity results in retention of a bound disc structure that does not instantly accrete onto the star.[43] For a given fluid element, we may then expect accretion onto the star to be mediated via a protoplanetary disc.[88, 66]
We now aim to generate a simplified ‘population synthesis’ of discs shaped by BHL accretion. To do so, we wish to initially remain agnostic as to the driver of stellar accretion.[89, 90] We therefore consider a stellar accretion rate $\dot{M}_{\mathrm{acc}}=M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$, for disc mass $M_{\mathrm{disc}}$ and fixed accretion time-scale $\tau_{\mathrm{acc}}$. For each of the discs in our sample discussed, we solve the initial value problem:

### v2

@@ 2.2 Properties of local star forming regions
In the first instance, we wish to quantify the properties of GMCs that undergo gravitational collapse to form stars. This can be reframed in terms of quantifying the maximum turbulent size-scale on which a given volume of space in the mid-plane of the galaxy is bound at any given time. All scales smaller than this size-scale will also be bound to to this cloud, and therefore this can be considered a unit of star formation, or a star forming region. This problem can be understood as a barrier-crossing problem arising from a random walk through Fourier space, which can be solved using excursion set theory as has been successfully applied to dark matter halos (Bond et al., 1991).
@@
We follow Hopkins (2012) who applied the excursion set formalism in the context of GMCs in the galactic disc. The approach can be summarised as follows, but we refer the interested reader to that work for greater detail. Most generally, it is clear that the density of gas on any given scale can be described as the sum of the contributions to this density at all larger scales. Supersonic turbulence yields a log-normal probability distribution in volume space for the gas density, with a dispersion that is dependent on the local Mach number (Vazquez-Semadeni, 1994). Therefore, dispersion in the contribution of any range of scales to the deviation from the median depends on the turbulent velocity on that scale. In a supersonic, rapidly cooling medium, the cascade of energy from large scales to small scales produces an energy spectrum with power-law exponent $p=2$, or turbulent velocity that scales with the square-root of the size-scale (Burgers, 1948). The critical density required for collapse at any intermediate size-scale is then related to the effective pressure support afforded to the medium from the turbulent motions on that scale, via the dispersion relation for a disc of a finite thickness (Begelman & Shlosman, 2009). On the largest scales, this pressure support effectively comes from galactic shear, while on the smallest scales, turbulent fragmentation can be suppressed by the thermal pressure (Jeans, 1902). In principle, the influence of magnetic fields can also be accounted for in this context, but Hopkins (2012) notes that this modification practically amounts to changing the sound speed, with only minor quantitative consequences for our model. In the context of the BHL accretion rate that we eventually compute, modifications of only a factor of order unity are expected due to local magnetic fields (Lee et al., 2014; Burleigh et al., 2017). Given the turbulent energy spectrum, a ‘trajectory’ in logarithmic density space can be instantaneously defined as the cumulative sum of perturbations from the largest to smallest scales. For any trajectory, a gravitationally unstable GMC, or star forming region, is defined on the largest length scale which is unstable (if any), all smaller length scales being subsumed within the larger volume.
@@
In our case, we follow similar parameter choices to those adopted by the fiducial case of Hopkins (2012). The energy spectrum normalisation is defined such that the velocity dispersion at the scale height $h$ of the gaseous galactic disc $\sigma_{\rm g}(h)=h\Omega=6$ km s-1, where $\Omega=2.6\times 10^{-2}$ Myr-1. We assume a surface density of the galactic disc in the solar neighbourhood $\Sigma=12\,M_{\odot}$ pc-2, yielding a Toomre $Q=1.33$. We always assume a sound speed $c_{\mathrm{s}}=0.2$ km s-1 and define a grid with a maximum scale $R_{\mathrm{max}}=5h$ down to a minimum scale $R_{\mathrm{min}}<0.01$ pc, each grid cell being $95$ percent of the spatial scale of the previous cell.
@@
where $t_{\rm cross}$ is the crossing time-scale and $\eta\approx 1$ (Pan & Scannapieco, 2010). An appropriate time-scale ${\tau}_{\mathrm{t,0}}$ over which the trajectory ‘resets’ corresponds to the turbulent time-scale on the scales that dominate density variations. We therefore estimate ${\tau}_{\mathrm{t,0}}\approx\tau_{\mathrm{t}}(h)$. We can then estimate the differential number of star forming region younger than a certain age $\Delta\tau$ within a given volume $\Delta V$:
@@
where $N_{\mathrm{coll}}$ is the number of unstable trajectories from a sample of $N_{\mathrm{traj}}$. We include a factor $1/2^{3}$ since given a region of radius $R_{\mathrm{SFR}}$, any other region with $R_{\mathrm{SFR}}$ is forbidden inside a distance of $2R_{\mathrm{SFR}}$ (such a region would merge with the original region). In the averaging for equation 2 we have ignored that similar mass unstable regions may be spatially correlated, since the probability of collapse on a given spatial scale is dependent on the density contribution from larger scales. However, Hopkins (2012) find that this bias is negligible for the low mass star forming regions, with $R_{\mathrm{SFR}}\ll h$. The shape of the mass function for the local, low mass GMCs relevant in the context of this work should therefore be correct. Clearly, star forming regions should still preferentially be found where the gas density is highest, we assume that using the average surface density for a sufficient $R_{\mathrm{obs}}\gtrsim h$.
We show the outcome of this calculation for $10^{5}$ independent trajectories in Figure 1. We plot the number of star forming regions above a given mass $M$ within a spherical volume with radius $R_{\mathrm{obs}}=200$ pc ($\sim h$) younger than $10$ Myr. This is approximately the radius within which the majority of protoplanetary disc statistics have been compiled (for a review, see Manara et al., 2023) and the dispersal time-scale for protoplanetary discs (Haisch et al., 2001; Pfalzner et al., 2022). We define the mass threshold below which we expect $>10$ star forming regions to be the threshold for being well sampled. This yields a mass limit $M_{\mathrm{local}}\approx 500\,M_{\odot}$. Note, our definition of a star forming region is a region that is collapsing to form stars. This is in contrast to a broader complex that may not be collapsing globally. As an illustrative example, the Taurus-Auriga association has a total mass much larger than $500\,M_{\odot}$ (Goldsmith et al., 2008). However, the complex as a whole is not collapsing to form stars. Instead collapse and star formation has occurred in several smaller sub-regions. It is these regions, or ‘NESTS’ (Joncour et al., 2018), that are the individual star forming regions in the sense appropriate to this work. In Taurus, these regions are typically a few tens of solar masses. The older region Upper Scorpius has a total stellar mass $\sim 1400\,M_{\odot}$ (Preibisch et al., 2002), but is also composed of several individual populations (Miret-Roig et al., 2022). The nearest individual star forming region with total mass of several $10^{3}\,M_{\odot}$ is the Orion Nebula cluster at a distance of $\sim 400$ pc (Hillenbrand & Hartmann, 1998; Reid et al., 2009). We conclude that, although our approach is approximate in that it neglects some physical processes that may influence local star formation, such as possible nearby supernovae (Zucker et al., 2022), the upper mass limit we infer is reasonable. Our subsequent results do not depend strongly on this limit, as long as we exclude much more massive regions that may be dispersed prematurely by stellar feedback.
@@
2.3 Collapse of the giant molecular cloud
@@
In the early stage after a star forms, it occupies the local star forming region that is undergoing gravitational collapse. Since we are interested in the local density and velocity evolution throughout a star’s lifetime, we require a simple model to follow the collapse of the GMC and formation of stars. Once the GMC becomes gravitationally unstable, it collapses on a free-fall time-scale. It does not do so monolithically, but because smaller sub-regions of the cloud may be higher density (and therefore have shorter free-fall time-scales), they fragment first and continue to accrete from their parent structures (Hoyle, 1953). Simultaneously, the cloud is exposed to heating by feedback processes from the stars as they form. A wide range of analytic and semi-analytic approaches have been developed to describe this process (Vázquez-Semadeni et al., 2019). We are helped in this context in our interest in low mass star forming regions that are locally well sampled. Since we do not expect many massive stars or supernovae among these local regions, we will hereafter ignore feedback processes and exclusively consider gravitational collapse.
@@
We adopt the approach of Girichidis et al. (2014) in modelling the rate of star formation during the collapse of the GMC. The approach of those authors is attractive for its simplicity, while also retaining a self-consistent treatment of the gas density distribution as it evolves under gravity. We briefly review the approach here, referring the reader to the original work for details.
@@
We consider a spherical region undergoing free-fall collapse from rest, which may be contained within a larger gravitationally unstable cloud. Girichidis et al. (2014) write an invertible, analytic approximation for the density of such a region as a function of time. Given an analytically integrable initial density distribution, there is a corresponding analytic expression for the fraction of an initial gas reservoir that collapses beyond any given threshold as a function of time. The initial density distribution in our collapsing region is log-normal, with a dispersion determined by the Mach number on the scale of collapse. We choose the critical density to be such that the local free-fall time is shorter than $0.1$ Myr. Given the accelerating nature of gravitational collapse, the exact threshold time is irrelevant so long as it is substantially shorter than the time-scale for the global collapse of the star forming region. The star formation rate is defined to be the rate at which material crosses this threshold density. We assume the cloud is instantly dispersed on cessation of the star formation phase, which continues until a star formation efficiency $\epsilon=0.5$ is reached, this value being commensurate with simulations of stellar core formation (Matzner & McKee, 2000). We expect stellar feedback, which we do not consider and may shut off star formation in massive clouds, to have a negligible influence on the low mass star forming regions that are our focus. High star formation efficiencies similar to those we assume are also reached for more sophisticated models of star formation in a collapsing cloud for low mass clouds (Vázquez-Semadeni et al., 2019).
2.4 Stellar dynamics in the bound cloud
@@
Having defined a local star formation rate, we can draw formation times for an initial stellar population. Since these stars are initially bound to the cloud from which they form, we integrate the orbit of the star within the gravitational potential of the cloud, ignoring interactions with other stars. For convenience we will consider a Plummer potential with a core radius $a=R_{\mathrm{SFR}}/2.6$, such that the outer radius of the star forming region is twice the half-mass radius. Stars are drawn from within the core radius, and we draw an isotropic velocity distribution with a dispersion assuming instantaneous virial equilibrium with the cloud. To track local density and relative velocity of the gas, we assume a static, smooth Plummer density profile (as for Throop & Bally, 2008). We then evolve the star under the potential, varying the core radius as the spherical star forming region contracts until the final dispersal of the cloud (when $M_{*}=\epsilon M_{\mathrm{tot}}$).
2.5 Evolution of the turbulent medium
@@
2.6 Accretion rate
Figure 2: The relative velocity (top) and density (middle) of the local gas, for a subset of six stars in our sample of 500 trajectories. The BHL accretion rate (bottom) is normalised to be independent of mass $m_{*}$, by which lines are coloured. The gas density and velocity are shown at the accretion scale (see text for details). In the middle panel, the horizontal dotted lines are the mean gas density for some high density regions in Taurus (Goldsmith et al., 2008). In the bottom panel, the red solid (dashed) line follows the median (mean) normalised BHL accretion rate for the entire sample, and a normalised two-dimensional histogram.
BHL accretion has been the topic of many works (for a review, see Edgar, 2004). In the purest case, it is the gravitational focusing of the test particles (representing the ISM), resulting in a stagnation point in the wake of the star within which material is accreted. The instantaneous BHL radius within which material is captured is then defined:
@@ 2.6 Accretion rate
Streamlines with an impact parameter at the BHL radius $R_{\mathrm{BHL}}$ hit a stagnation point along the axis coincident with the relative upstream gas velocity vector. In the ideal case, in an homogeneous medium, they then accrete directly onto the star. However, in some circumstances the radius within which capture can occur can be smaller than the BHL radius. Firstly, clearly the accretion radius cannot exceed the turbulent size scale under consideration, $R_{\mathrm{t}}$. In addition, if the tidal radius $R_{\mathrm{T}}<R_{\mathrm{BHL}}$, then this radius can also limit the material that can be captured by the star. If a star is bound to a mass $M_{\mathrm{b}}$ at a radius $r_{\mathrm{b}}$, then:
@@
Initially, this radius is determined by the collapsing cloud, where $r_{\mathrm{b}}$ is the instantaneous location of the star with respect to the centre of the region. Once the cloud has dispersed, we take the galactic values $r_{\mathrm{b}}=8$ kpc and $M_{\mathrm{b}}=10^{12}\,M_{\odot}$. Given these considerations, the accretion radius can be defined:
@@
Material within the cross-section carved out by this radius is captured by the star. At any turbulent scale, we can estimate the instantaneous rate at which a star captures material in such a scenario (Bondi, 1952; Shima et al., 1985):
@@
We show the outcome for the evolution of the local density, relative velocity and BHL accretion rate (normalised to show the accretion rate assuming a solar mass star) for a subset of six stars in Figure 2. Given that density in particular can vary over several orders of magnitude, we also draw an approximate comparison to the typical densities in observed star forming regions. To do so, we adopt averaged density estimates from Table 4 of Goldsmith et al. (2008) for a number of high density areas within the Taurus star forming complex. We convert the mass estimates $M_{\mathrm{enc}}$ within the quoted surface area $A_{\mathrm{surf}}$ to a density estimate by approximating spherical geometry, such that the volume averaged density is:
@@
The resultant BHL accretion rates, normalised for the a solar mass star, are shown in Figure 2, for a subset of six stars in our sample. The phase in which the star remains within the bound gaseous reservoir typically persists for $\sim 1$ Myr. During this phase, local low velocity and high density gas results in an early phase of enhanced BHL accretion at a median rate $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-8}\,M_{\odot}$ yr-1 (Figure 2, bottom). This median rate then quickly drops to $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-10}{-}10^{-11}\,M_{\odot}$ yr-1 in the age range $1{-}3$ Myr for a solar mass star. This is considerably lower than the typical rate of accretion onto the central star, $\dot{M}_{\mathrm{acc}}\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 (Manara et al., 2023). One might therefore assume that BHL accretion no longer plays an important role for disc evolution. However, in a turbulent medium, vorticity of the in-falling material means that accretion is mediated via a (protoplanetary) disc (Krumholz et al., 2006; Kuffmeier et al., 2020). In this case, the change in accretion rate can be averaged over the time-scale $\tau_{\mathrm{acc}}$ at which angular momentum is transported within the disc. Empirically, this time-scale can be a considerable fraction of the disc lifetime (Almendros-Abad et al., 2024). In this case the median underestimates the true contribution of BHL accretion rate onto the disc to the accretion rate onto the star. By contrast, the mean accretion rate when averaged over all stars in our model is $\langle\dot{M}_{\mathrm{BHL}}\rangle\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 at $1-3$ Myr, as shown as the dashed red line in Figure 2. However, the mean overestimates the typical BHL accretion rate contribution for individual discs, since it is dominated by a small number of rapid accretors. Interpreting the contribution from BHL accretion to the stellar accretion rate for typical star-disc systems is therefore not a simple task. Compounding this problem further, discs that survive to comprise the observed sample may also preferentially be those which have been replenished by the ISM. Interpreting the role of BHL accretion for disc evolution therefore requires further calculation.
3 Protoplanetary disc properties
3.1 Simple disc model
In an homogeneous medium the velocities of steamlines are zero at the stagnation point. Captured material therefore has zero angular momentum and accretes directly onto the star. However, in an inhomogeneous medium, vorticity results in retention of a bound disc structure that does not instantly accrete onto the star (Krumholz et al., 2006). For a given fluid element, we may then expect accretion onto the star to be mediated via a protoplanetary disc (Moeckel & Throop, 2009; Wijnen et al., 2017).
We now aim to generate a simplified ‘population synthesis’ of discs shaped by BHL accretion. To do so, we wish to initially remain agnostic as to the driver of stellar accretion (Lynden-Bell & Pringle, 1974; Tabone et al., 2022). We therefore consider a stellar accretion rate $\dot{M}_{\mathrm{acc}}=M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$, for disc mass $M_{\mathrm{disc}}$ and fixed accretion time-scale $\tau_{\mathrm{acc}}$. For each of the discs in our sample discussed, we solve the initial value problem: